HUBBLE SPACE TELESCOPE DAILY REPORT #840 PERIOD COVERED: 0800 EST Daily Status Report as of 069/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.2 Completed PC 3769 (Structure of Inner Coma of Comet Chiron) The Planetary Camera was used with filter F555W to make four observations of the comet/object 2060 Chiron. This is to find its outer edge (i.e., the exopause) of the gravitationally bound dust cloud around the nucleus (i.e., the inner coma), and also to look for indications of active regions in the inner coma. All four observations were received and displayed in the OSS. They were reported to look okay. 1.3 Completed WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) Both the wide Field Camera and the Planetary Camera were used with filter F555W to make internal flat observations to check for new forms of contamination (i.e. daisies and measles). Both internal flat observations were reported to look normal. 1.4 Completed HSP/POL 3557 (Polarimetric Observations of Crab Pulsar in the UV) Because of the fact that the High speed Photometer had been placed in safe mode (064/2344Z), the four planned observations of the Crab Nebula Pulsar did not occur. (See Operations Note 2.4) 1.5 Completed WFC 1306 (GTO Low Latitude Parallel Program) The Wide Field Camera was used to make two low latitude observations taken in parallel with the prime science observations. Both observations were received and displayed. They both were reported to look okay. 1.6 Completed S/C/FOS/BL/RD 4098 (FOS Discriminator Test) The Faint Object Spectrograph was used to make one spectral calibration observation of determine to optimal diode discriminator settings. The observations was received and displayed. It was reported to look normal. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 11 Successful : 11 Losses of Lock : 1 (LOL AT 069/0129Z, HSTAR 3909) 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Significant HSTARs: None 2.4 Operations Note: 2.4 Operations Note: On Monday (Day 067) while the High Speed Photometer was being used to make polarimetric observations of the Crab nebula pulsar the Science Data Formatter (SDF) unit locked up. The SDF lock up occurred because of improper handshaking between the NSSC-I and the SDF. The HSP was disabled to enable other SI observations to continue. Shortly afterwards, due to concerns over the vulnerability of the HSP to bright objects, the HSP was safed. The command plan is currently being modified to allow the HSP observations to resume as soon as is possible. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/BL 3538 ( UV spectra of Extended Emission Regions Around QSOs) 3.2 Complete Two Sets of S/C/FOS/BL 4099 (FOS focus, X-Pitch and Y-Pitch) 3.3 Complete FOC/96 4800 (Stellar content of Wolf-Rayet Galaxies) 3.4 Complete WFC 4658 (Determination of Extragalactic Distance Scale) 3.5 Complete GHRS/2 3626 (Ephemeral Determination fo wind Velocities and Density Laws) 3.6 Complete PC 4140 (WF/PC Calibration: Rapid Internal Monitor Part 2) 3.7 Complete Uplink Updated Ephemeris Tables, Per MOSES 2871 3.8 Complete GHRS/2 3496 (Ultraviolet Atlas of Sirius) 3.9 Initiate FOC/96 3591 (Massive Stars in Starburst galaxies - FOC images) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-